### v1

@@ Abstract
Despite the similar structures of human faces, existing face alignment methods cannot learn unified knowledge from multiple datasets with different landmark annotations. The limited training samples in a single dataset commonly result in fragile robustness in this field. To mitigate knowledge discrepancies among different datasets and train a task-agnostic unified face alignment (TUFA) framework, this paper presents a strategy to unify knowledge from multiple datasets. Specifically, we calculate a mean face shape for each dataset. To explicitly align these mean shapes on an interpretable plane based on their semantics, each shape is then incorporated with a group of semantic alignment embeddings. The 2D coordinates of these aligned shapes can be viewed as the anchors of the plane. By encoding them into structure prompts and further regressing the corresponding facial landmarks using image features, a mapping from the plane to the target faces is finally established, which unifies the learning target of different datasets. Consequently, multiple datasets can be utilized to boost the generalization ability of the model. The successful mitigation of discrepancies also enhances the efficiency of knowledge transferring to a novel dataset, significantly boosts the performance of few-shot face alignment. Additionally, the interpretable plane endows TUFA with a task-agnostic characteristic, enabling it to locate landmarks unseen during training in a zero-shot manner. Extensive experiments are carried on seven benchmarks and the results demonstrate an impressive improvement in face alignment brought by knowledge discrepancies mitigation.
@@ 1 Introduction
Moreover, by mitigating the knowledge discrepancies using the proposed strategy, the learned knowledge is extended from target landmarks to face structure. The learned structure knowledge and the proposed structure prompts significantly improve the knowledge transfer efficiency when apply a pretrained model to a group newly-defined landmarks. As a result, the performance of few-shot learning can be effectively improved. Moreover, explicitly representing face on the interpretable plane also makes the learned model task-agnostic. By encoding the coordinates on the interpretable plane into structure prompts, TUFA also achieves zero-shot face alignment for the first time, predicting not only the pre-defined landmarks, but also those landmarks unseen during training, as shown in Fig. 2.
@@ 2.1 Many-shot Face Alignment
As a very pivotal technology in computer vision, many-shot face alignment has undergone over two decades of intensive research and exploration. At the very early stage, constrained local model (CLM) [11], active shape model (ASM) [10], active appearance model (AAM) [9, 1, 37] dominated this task. Their core idea is to align a statistical mean shape to a new face and that is why this task is named as face alignment rather than facial landmark detection. To overcome the fragile robustness and generalization, cascaded shape regression (CSR) methods [6, 67, 19, 79, 18] are proposed to achieve face alignment with multi-steps. Lately, the development of CNN and vision transformer (ViT) [16] significantly boost the performance of face alignment. Recent many-shot face alignment methods can be roughly divided into two categories: heatmap regression methods [58, 43, 69, 53, 15, 65, 5, 12, 59, 57, 8, 81, 50, 72, 27, 29, 32, 76, 31, 24, 75] and coordinate regression methods [55, 66, 56, 41, 61, 28, 60, 17, 62, 36, 14, 78, 34, 35, 77, 63, 64]. They set each landmark as an independent regression target and their learned models are task-specific.
@@ 3.3 Loss Function
Figure 5: (a) the red points indicate the randomly generated scratch shape, and the blue points represent the mean shape of 98 landmarks, which serves as anchors during training. (b) the red points represent the zero-shot face alignment results under various conditions.
@@ 4.4 Comparisons in Within-dataset Validation
WFLW: we report the inter-ocular NME, parameter size and flops of TUFA and other state-of-the-art methods in Table 1. FR0.1 and AUC0.1 are tabulated in Table 2. The TUFA learned from multiple datasets yields best performance in terms of NME on the full set and all subsets. The unique semantic alignment enables TUFA to maximize the utilization of extra face images. Even though TUFA employs fewer extra samples compared to other methods trained with multiple datasets, we still observe a significant improvement of 7.09% in NME compared to the TUFA trained with a single dataset. This demonstrates that TUFA can maximize the utilization of extra training samples.
@@
TUFA (ViT-S/16) 2.68 4.58 3.05
DSLPT (ViT-S/8) 2.59 4.45 2.95
Considering efficiency, we have also implemented a TUFA with ViT-S/16. Despite having comparable flops to other methods, it yields the second best performance in NME and FR0.1. Moreover, it does not require any post-processing as it regresses the coordinates of landmarks. Therefore, this TUFA runs faster than heatmap regression methods even though they have similar flops. The entire training process of this TUFA can be completed within 5 hours on a single A40 GPU.
@@ 4.5 Comparisons in Cross-dataset Validation
[…]
Table 9: Performance comparison in inter-ocular NME (%$\downarrow$ with reduced training set on WFLW full set. Key: [ Best]
@@
Table 10: Performance of TUFA in inter-ocular NME ($N_{\rm pre}$=10) on CelebA under the setting of zero-shot face alignment. The evaluation results of state-of-the-art self-supervised and unsupervised methods also reported. Key: [ Best]
@@
4.6 Few-shot Face Alignment
We pretrain TUFA (ViT-S/8) on COFW, 300W and MERL-RAV (29 & 68 landmarks) and implement few-shot face alignment on WFLW (98 landmarks) as described in Section 3.4. Note that there are large differences between the semantic definitions of 29 & 68 landmarks and 98 landmarks. We randomly select a certain number of training samples in each experiment. With varying training set sizes, we repeat each experiment five times, and the mean NME with each training set size is reported in Table 9. With only 10 training samples, TUFA still demonstrates superior performance compared to other methods, outperforming 3FabRec and He et al. trained with 7500 annotated faces. When we fine-tune TUFA on WFLW with 100% of the training set, the NME on the testing set drops to as low as 3.99%. This performance is comparable to the NME of the TUFA when trained with all four datasets together. This result indicates that TUFA can effectively leverage the knowledge learned from pretraining. This is achieved because TUFA successfully unifies the learning objectives of different datasets and mitigates their knowledge discrepancies, rather than treating landmark regression as an independent target. Because of this unique property, TUFA can be transferred to a group of newly defined landmarks efficiently, even with a very limited number of annotated samples. 3FabRec, Autolink and He et al. set image reconstruction as a pre-text task to encourage model to perform better with very limited training images. However, He et al. also found that this pre-text task may also lead to performance degradation when the number of training samples is sufficient. The main reason is that the learning objectives of image reconstruction and face alignment are not aligned. This large knowledge discrepancy significantly degrades performance.
4.7 Zero-shot Face Alignment
@@
Table 11: Performance of TUFA in inter-ocular NME (%$\downarrow$ on WFLW under the setting of zero-shot face alignment. The evaluation results of some fully supervised methods before 2018 are also reported. Key: [ Best]
@@
Figure 8: Zero-shot alignment results predicted by TUFA on WFLW. The red points and green points represent predicted landmarks and ground truth respectively.
Table 12: The influence of different numbers of datasets used in training. The inter-ocular NME on 300W, COFW and WFLW, and the NMEbox on MERL-RAV are reported. Key: [ Best]
@@
Table 15: The influence of masking ratio on WFLW. The inter-ocular NME is reported. Key: [ Best]
@@ 4.8 Ablation Studies
Influence of the masking ratio: to further explore the influence of masking ratio, we train TUFA with different masking ratios on WFLW. The results are reported in Table 15. As the masking ratio increases, the performance of TUFA is improved. The primary reason is that randomly masking a certain ration of landmarks encourages TUFA to learn the mapping using the relationship between distant landmarks. This relationship ensures a more coherent mapping between the 2D plane and target faces, while the landmarks that are close together often introduce bias into the learned mapping. The results show TUFA works well within the range of $[50\%,85\%]$. If the masking ratio exceeds 85%, the number of remaining landmarks, which serves as anchors during training, is insufficient to construct a mapping from a 2D plane to a highly complex face surface. As a result, the performance of TUFA degrades significantly.
@@ 5 Conclusion
In this paper, we introduce a novel strategy to mitigate knowledge discrepancies among multiple datasets and further achieve task-agnostic, unified face alignment. Compared to existing face alignment methods, the proposed method represents a significant breakthrough in three critical aspects: 1) TUFA successfully mitigates knowledge discrepancies among multiple datasets by aligning their landmark semantics on an interpretable plane and further unifying their learning target to map the plane to target faces. 2) the unified learning target also enables the learned knowledge to be easily transferred to a new set of defined landmarks. Consequently, TUFA significantly boosts the performance of few-shot face alignment. 3) By mitigating the discrepancies, the learned knowledge is extended from target landmarks to face structure. We employ structure prompts to further bridge the gap between seen and unseen landmarks, achieving zero-shot face alignment for the first time. Overall, we believe the unique properties of TUFA can propel face alignment to a new stage and provide a competitive baseline for future works.

### v2

@@ Abstract
Despite the similar structures of human faces, existing face alignment methods cannot learn unified knowledge from multiple datasets with different landmark annotations. The limited training samples in a single dataset commonly result in fragile robustness in this field. To mitigate knowledge discrepancies among different datasets and train a task-agnostic unified face alignment (TUFA) framework, this paper presents a strategy to unify knowledge from multiple datasets. Specifically, we calculate a mean face shape for each dataset. To explicitly align these mean shapes on an interpretable plane based on their semantics, each shape is then incorporated with a group of semantic alignment embeddings. The 2D coordinates of these aligned shapes can be viewed as the anchors of the plane. By encoding them into structure prompts and further regressing the corresponding facial landmarks using image features, a mapping from the plane to the target faces is finally established, which unifies the learning target of different datasets. Consequently, multiple datasets can be utilized to boost the generalization ability of the model. The successful mitigation of discrepancies also enhances the efficiency of knowledge transferring to a novel dataset, significantly boosts the performance of few-shot face alignment. Additionally, the interpretable plane endows TUFA with a task-agnostic characteristic, enabling it to locate landmarks unseen during training in a zero-shot manner. Extensive experiments are carried on seven benchmarks and the results demonstrate an impressive improvement in face alignment brought by knowledge discrepancies mitigation. The code is available at https://github.com/Jiahao-UTS/TUFA.
@@ 1 Introduction
Moreover, previous face-alignment methods based on heatmap regression and coordinate regression treat each landmark regression as a separate task, resulting in task-specific knowledge. Although human faces share a similar structure, it remains challenging to extend this knowledge to unseen facial landmarks. Nevertheless, by mitigating the knowledge discrepancies using the proposed strategy, the learned knowledge is not only limited to the labeled landmarks in training but also includes task-agnostic knowledge: the entire face structure. The learned task-agnostic knowledge can be efficiently transferred to a group of newly defined landmarks using only a few labeled samples, which significantly boosts the performance of few-shot face alignment. Moreover, by explicitly representing the face on an interpretable plane and encoding its coordinates into structure prompts, this knowledge can also be used to localize unseen facial landmarks without fine-tuning. Therefore, TUFA also achieves zero-shot face alignment for the first time, predicting not only the pre-defined landmarks, but also those landmarks unseen during training, as shown in Fig. 2.
@@ 2.1 Many-shot Face Alignment
As a very pivotal technology in computer vision, many-shot face alignment has undergone over two decades of intensive research and exploration. At the very early stage, constrained local model (CLM) [11], active shape model (ASM) [10], active appearance model (AAM) [9, 1, 34] dominated this task. Their core idea is to align a statistical mean shape to a new face and that is why this task is named as face alignment rather than facial landmark detection. To overcome the fragile robustness and generalization, cascaded shape regression (CSR) methods [6, 62, 15, 73, 14] are proposed to achieve face alignment with multi-steps. Lately, the development of CNN and vision transformer (ViT) [13] significantly boost the performance of face alignment. Recent many-shot face alignment methods can be roughly divided into two categories: heatmap regression methods and coordinate regression methods. They set each landmark as an independent regression target and their learned models are task-specific.
@@ 3.3 Loss Function
Figure 5: (a) the red points are used to generate face structure prompts for zero-shot face alignment. The blue points represent the mean shape of 98 landmarks, which serves as anchors during training. (b) the red points represent the zero-shot face alignment results under various conditions.
@@ 4.4 Comparisons in Within-dataset Validation
WFLW: we report the inter-ocular NME, parameter size and flops of TUFA and other state-of-the-art methods in Table 1. FR0.1 and AUC0.1 are tabulated in Table 2. Without increasing any computational complexity, the ability of TUFA o mitigate knowledge discrepancies across multiple datasets maximizes the utilization of extra training samples. We observe a significant improvement of 7.09% in NME on the full set compared to TUFA (ViT-S/8) trained on a single dataset, even though TUFA uses fewer extra samples than other methods trained on multiple datasets.
Considering efficiency, we also implemented a TUFA with a lighter backbone, ViT-S/16. Similarly, training on multiple datasets achieves an impressive improvement of 8.68% in NME on the full set compared to single dataset training. Despite having much lower computational complexity (flops) than most heatmap regression methods, it still yields the second best performance in NME and FR0.1. It also outperforms DSLPT in NME on the full set and all subsets with comparable flops. Moreover, the entire training process for this TUFA can be completed within five hours on a single A40 GPU, demonstrating the high efficiency of our method.
@@
TUFA (ViT-S/8) 2.59 4.45 2.95
@@ 4.5 Comparisons in Cross-dataset Validation
Table 9: Performance comparison in inter-ocular NME (%$\downarrow$ with reduced training set on WFLW full set. Key: [ Best, Second Best]
@@
4.6 Few-shot Face Alignment
We pretrain TUFA (ViT-S/8) on COFW, 300W and MERL-RAV (29 & 68 landmarks) and implement few-shot face alignment on WFLW (98 landmarks) as described in Section 3.4. Note that there are large differences between the semantic definitions of 29 & 68 landmarks and 98 landmarks. We randomly select a certain number of training samples in each experiment. We repeat each experiment five times and report the mean NME with each training set size in Table 9. With only 10 training samples, TUFA still demonstrates superior performance compared to other methods, outperforming 3FabRec and He et al. trained with 7500 annotated faces. When we fine-tune TUFA on WFLW using 100% of the original training set (7500 samples), the NME on the testing set improves to 3.99%. This performance is comparable to the NME of the TUFA when trained with all four datasets together. This result indicates that TUFA can effectively leverage the knowledge learned from pretraining. This is achieved because TUFA successfully unifies the learning objectives of different datasets and mitigates their knowledge discrepancies, rather than treating landmark regression as an independent target. Because of this unique property, TUFA can be transferred to a group of newly defined landmarks efficiently, even with a very limited number of annotated samples. 3FabRec, Autolink and He et al. set image reconstruction as a pre-text task to encourage model to perform better with very limited training images. However, He et al. also found that this pre-text task may also lead to performance degradation when the number of training samples is sufficient. The main reason is that the learning objectives of image reconstruction and face alignment are not aligned. This large knowledge discrepancy significantly degrades performance.
4.7 Zero-shot Face Alignment
Table 10: Performance of TUFA in inter-ocular NME ($N_{\rm pre}$=10) on CelebA under the setting of zero-shot face alignment. The evaluation results of state-of-the-art self-supervised and unsupervised methods also reported. Key: [ Best, Second Best]
@@ 4.7 Zero-shot Face Alignment
Table 11: Performance of TUFA in inter-ocular NME (%$\downarrow$ on WFLW under the setting of zero-shot face alignment. The evaluation results of some fully supervised methods before 2018 are also reported. Key: [ Best, Second Best]
@@
Table 12: The influence of different numbers of datasets used in training. The inter-ocular NME on 300W, COFW and WFLW, and the NMEbox on MERL-RAV are reported. Key: [ Best, Second Best]
@@
Table 15: The influence of masking ratio on WFLW. The inter-ocular NME, FR0.1, AUC0.1 and the computational complexity of Decoder (Flops) are reported. Additionally, the memory usage and training speed on a single A40 are also provided (with the batch size set to 16). Key: [ Best, Second Best]
@@
Decoder Flops (G) $\downarrow$ 3.537 3.089 2.663 2.226 2.139 2.053 1.967
Memory usage (MB)$\downarrow$ 14,004 13,792 13,608 13,502 13,490 13,436 13,368
Training Speed (samples/sec) $\uparrow$ 45.8 46.4 46.8 47.2 47.3 47.3 47.5
Figure 8: Zero-shot alignment results predicted by TUFA on WFLW. The green points and red points represent the predicted landmarks and ground truth respectively.
@@ 4.8 Ablation Studies
Influence of the masking ratio: to further explore the influence of masking ratio, we train TUFA (ViT-S/8) with different masking ratios on WFLW. The results are reported in Table 15. As the masking ratio increases, the performance of TUFA is improved. The primary reason is that randomly masking a certain ration of landmarks encourages TUFA to learn the mapping using the relationship between distant landmarks. This relationship ensures a more coherent mapping between the 2D plane and target faces, while the landmarks that are close together often introduce bias into the learned mapping. Moreover, increasing the masking ratio also significantly reduces the computational complexity of the TUFA decoder. When the ratio is set to 75%, the computational complexity of TUFA decoder is only 2/3 of that at a 0% masking ration. Although TUFA decoder only accounts for just about 10% of the total computational complexity compared to TUFA encoder (31.572G flops), we can still observe a notable reduction in memory usage and an acceleration in training speed brought by the masking training strategy.
The results show TUFA works well within the range of $[50\%,85\%]$. If the masking ratio exceeds 85%, the number of remaining landmarks, which serves as anchors during training, is insufficient to construct a mapping from a 2D plane to a highly complex face surface. As a result, the performance of TUFA degrades significantly. After comprehensively considering both performance and computational complexity, we set the masking ratio to 0.75 in our model.
@@
4.9 Limitations
TUFA can only generalize to unseen landmarks within the face area. If TUFA uses a set of structural prompts generated from points outside the face area on the interpretable plane, it still localize a set of nearest landmarks to these points inside the face area. The main reason is that the training datasets only contain information about facial structure. By using a dataset for dense human pose estimation [17], it becomes possible to locate unseen landmarks on human bodies in a zero-shot manner.
@@ 5 Conclusion
In this paper, we introduce a novel strategy to mitigate knowledge discrepancies among multiple datasets and further achieve task-agnostic, unified face alignment. Compared to existing face alignment methods, the proposed method represents a significant breakthrough in three critical aspects: 1) TUFA successfully mitigates knowledge discrepancies among multiple datasets by aligning their landmark semantics on an interpretable plane and further unifying their learning target to map the plane to target faces. 2) the unified learning target also enables the learned knowledge to be easily transferred to a new set of defined landmarks. Consequently, TUFA significantly boosts the performance of few-shot face alignment. 3) By mitigating the discrepancies, the learned knowledge is extended from target landmarks to face structure. As a result, we can employ structure prompts to further bridge the gap between seen and unseen landmarks, achieving zero-shot face alignment for the first time.
Although the gap between seen and unseen landmarks has been successfully addressed, the gap between different domains, such as varying lighting conditions, has not been explicitly considered in TUFA. In the future, we will explore the domain gap among facial landmarks to develop a more universal face alignment approach.
Declarations
Data Availability Statement: the datasets used in our paper are available in: